If you borrow $2000 and charged 7.0% interest compounded monthly and pay the principal interest in full after 72 weeks. a. how much money will you owe? b. how much interests will they make off the loan! can you walk me through the problem?
Answer by mananth(16946)

Principal P = 2000
Amount= A
years=n 6.00
compounded 12 times a year t
Rate = 7.00 0.01
Amount = P*((n+r)/n)^n*t

Amount =2000*(1 +0.01)^ 6*12
Amount =2000*(1 +0.01)^72
2000 *( 1.01 )^ 72
Amount = 3040.21

Interest = 3040 - 2000 = $1,040